# Initialization of Monocular Visual Navigation for Autonomous Agents Using Modified Structure from Small Motion

### v1

@@ IV CONCLUSIONS
In this work, we present a monocular initialization pipeline that enhances the 3-step SfSM approach to address visual estimation in weak-perspective projection, center-pointing trajectories, and presence of dominant planes for space vSLAM applications. We demonstrate the robustness of the proposed method in both the theoretical formulation and the experimental evaluation on realistic simulated image sets, and we prove the proposed method’s increased accuracy over the original 3-step SfSM method.
Future work for the proposed method should address the computation time and robustness of the initializer so that the approach may be implemented in monocular SLAM pipelines applied to more real-time, time-stringent applications. Moreover, variable pre-conditioning may be used to increase the rate of convergence of step 3 of the pipeline, addressing the vanishing Jacobian issue. Other possible routes of improvement include, but are not limited to, the use of simulated annealing [30] for better exploration of local minima, as well as convex relaxation methods [31] for an improved nonlinear program initialization.

### v2

@@ IV CONCLUSIONS
In this work, we present a monocular initialization pipeline that extends the Ha et al. three-step SfSM method [11] to address visual estimation in RSO inspection trajectories, which are characterized by weak-perspective projection, center-pointing trajectories, dominant planar geometry, and dynamic illumination conditions. We demonstrate improved accuracy and robustness over Ha’s approach and the USAC_FM_8PTS wide-baseline method when evaluating on realistic simulated image sets.
Future work for the proposed method should focus on improving the computation time and further robustness of the initializer, enabling its integration into monocular SLAM pipelines for real-time, time-critical applications. For instance, variable pre-conditioning may be used to increase the rate of convergence of step 3 of the pipeline, addressing the vanishing Jacobian issue. Other potential improvements include using simulated annealing [29] for better local minima exploration and applying convex relaxation methods [30] to improve nonlinear program initialization. Incorporating motion constraints for different RSO motion paradigms may also help further constrain the problem.